# SDC-Net: A Domain Adaptation Framework with Semantic-Dynamic Consistency for Cross-Subject EEG Emotion Recognition

### v1

@@ V-B Confusion matrices
In this section, we compare the performance of four different models: Da-CapsNet, PLMSDANet, LGDAAN-Net and our proposed model, SDC-Net, using confusion matrices (as shown in Fig. 2). The confusion matrices provide insight into how well each model classifies inputs into three categories: Negative, Neutral, and Positive. The diagonal elements represent correct predictions, while off-diagonal elements correspond to misclassifications. Our proposed SDC-Net model outperforms the other models, particularly in classifying the Neutral class with an accuracy of 92.45%, which is the highest among all models.
@@
Additionally, SDC-Net achieves 95.7% accuracy in the Positive class and 87.23% in the Negative class. The confusion between classes is minimal, with only 5.25% of neutral instances being misclassified as negative and 2.53% of positive instances being misclassified as negative. The low misclassification rates in SDC-Net suggest that our model effectively captures subtle differences between sentiment classes, particularly between negative and neutral sentiments, where other models faltered. It achieves the best accuracy in both the Neutral and Positive classes and exhibits significantly lower misclassification rates compared to the other models. This indicates that SDC-Net provides a more nuanced understanding of sentiment, making it particularly effective for tasks that require fine-grained sentiment analysis. In contrast, models like Da-CapsNet, despite excelling in classifying positive samples, demonstrate significant weaknesses in distinguishing negative and neutral sentiments, limiting their effectiveness in comprehensive sentiment classification tasks.
V-C Ablation Study
@@
To evaluate the contribution of each module in the proposed SDC-Net, we performed ablation experiments by systematically removing six key components. Table VI summarizes the performance under each setting in terms of average Pearson correlation coefficient (Pcc). First, domain alignment mechanisms were found to be essential. Removing the MMD component resulted in a significant drop in performance (86.82%), indicating its crucial role in reducing marginal distribution discrepancy between source and target domains. Similarly, excluding CMMD led to a moderate performance decline (89.91%), showing that conditional alignment further refines domain adaptation, though its impact is secondary to MMD. Second, the effect of DSCL was also prominent. Supervised similarity consistency on source domain contributes to extracting discriminative features from labeled source data, and its removal decreased the accuracy to 88.99%. More critically, excluding self-inferred similarity consistency leanring on target domain reduced performance to 86.99%, emphasizing its importance in modeling intra-class similarity in the unlabeled target domain, thus supporting better generalization. Third, SS-Mix based data augmentation improved generalization by synthesizing diverse EEG trials. Its removal caused a noticeable decrease in Pcc (87.99%), suggesting its role in mitigating overfitting and increasing sample diversity. Finally, the pseudo-confidence mechanism, which filters unreliable pseudo-pairs, slightly improved the mean accuracy and significantly reduced variance. Without this mechanism, the model still achieved 89.66%, but with increased performance fluctuation (standard deviation 7.18), indicating its stabilizing effect on target domain predictions.
In summary, MMD and target similarity consistency learning are the most influential modules in our framework. The remaining components also enhance the model’s robustness and consistency, contributing to the superior performance of SDC-Net across cross-subject EEG emotion recognition tasks.
TABLE VI: Performance of the SDC-Net model in the ablation study
Ablation Experiment Strategy Pcc (%)
without-SS-Mix 87.99 $\pm$ 5.77
@@
Figure 6: t-SNE visualization of source and target domain emotion representations across training stages.
V-D Visualization of Domain Alignment via t-SNE
To qualitatively assess the effectiveness of domain alignment, we visualized the feature distributions of source and target domains using t-SNE at different training stages (Fig. 6). In these plots, the source domain ($D_{s}$) is represented by circles and the target domain ($D_{t}$) by stars, with emotion categories indicated by color (blue: negative, red: neutral, green: positive). At the early training stage, features from $D_{s}$ and $D_{t}$ are poorly aligned. The target samples, especially for Subject 1 and Subject 2, are scattered and show significant overlap across emotion classes, indicating large domain discrepancy. Positive target samples (green stars) frequently overlap with other categories, suggesting the initial failure in feature alignment. As training progresses, domain alignment improves notably. For Subjects 2 and 3, the target domain emotion clusters begin to align with those of the source domain. Notably, negative and positive emotions become more separable, suggesting that MMD and CMMD modules effectively minimize marginal and conditional distribution gaps. By the final training stage, most emotion clusters exhibit clear separation in both domains. Subject 3 shows the most consistent alignment across categories, while Subject 1 retains partial overlap in neutral emotions. This aligns with known challenges in classifying neutral emotional states, which often lie closer to the decision boundaries due to their ambiguous EEG patterns.
Overall, the t-SNE visualizations provide intuitive support for the proposed framework’s ability to learn emotion-discriminative and domain-invariant features. Although some overlap remains—particularly for neutral states—the progressive clustering over epochs confirms the effectiveness of the domain adaptation strategy.
Figure 7: Hyperparameter Analysis of the Multi-Scale Gaussian Kernel in RKHS-based Distribution Alignment.
V-E Sensitivity Analysis for Hyperparameters
Figure 8: Hyperparameter Analysis of Similarity Consistency Thresholds.
Figure 9: Hyperparameter Analysis of Batch-Size.
@@ V-E Sensitivity Analysis for Hyperparameters
Figure 10: Topographic analysis of the mutual information between the EEG patterns and the model predictions.
V-E1 Gaussian Kernel Hyperparameters
To evaluate the effect of kernel configuration on model performance, we analyze combinations of kernel number ($L\in{1,3,5,7,10}$) and bandwidth multipliers ($\kappa\in{1,2,4,8}$). As shown in Fig. 7, increasing $L$ from 1 to 5 consistently improves classification accuracy. The highest performance is achieved at $L=5$, $\kappa=2.0$ ($91.85\%\pm 5.98\%$), balancing local and global distribution alignment. However, using too many kernels ($L>5$) introduces redundancy and noise, slightly degrading performance.
Moderate bandwidth scaling ($\kappa=2.0$) offers the best trade-off: small $\kappa$ may overfit local variations, while large $\kappa$ over-smooths inter-class differences. Overall, a small ensemble of Gaussian kernels with moderate scaling enables more effective representation alignment in RKHS.
V-E2 Effect of Similarity Consistency Thresholds
We further examine the sensitivity of the similarity consistency mechanism by varying the upper and lower thresholds $(\tau_{pu},\tau_{pl})$ used to define confident sample pairs in the target domain (see Fig. 8). The best performance ($91.85\%\pm 5.23\%$) is observed at $(\tau_{pu}=0.9,\tau_{pl}=0.5)$, indicating that enforcing constraints only on highly confident pairs improves stability and accuracy. Narrow threshold gaps (e.g., $(0.8,0.6)$) reduce the filtering of noisy relations, while looser settings (e.g., $\tau_{pl}=0.3$) introduce more false dissimilarities. These results suggest that carefully calibrated thresholds are critical for filtering uncertainty and ensuring stable pairwise supervision.
V-E3 Effect of Batch-Size
To investigate the influence of batch-size on model performance, we evaluate five different settings: 32, 64, 96, 128, and 256. As shown in Fig. 9, the model achieves the highest average accuracy when the batch-size is set to 32, with performance gradually degrading as the batch-size increases. Notably, although batch-size = 64 and 96 still maintain relatively competitive results, large batch sizes (128 and 256) exhibit increased performance variance and a downward trend in mean accuracy. This phenomenon can be attributed to the trade-off between gradient estimation stability and generalization ability. Smaller batch sizes provide noisier but more generalizable gradient estimates, whereas larger batches lead to more stable updates but risk overfitting to dominant features. In our case, the batch-size of 32 strikes a better balance between convergence stability and generalization, yielding an average accuracy of 90.35% and a maximum of 100%. These findings suggest that an appropriately small batch-size is more beneficial for emotion-related EEG tasks, potentially due to the high variability and subject-specific nature of EEG signals.
V-F Topographic Analysis of Important EEG Patterns
[…]
By computing the MI between each EEG feature and the predicted probabilities of the emotion categories, an initial MI $MI\in R^{3\times 5\times 2}$ tensor is derived, capturing the nonlinear statistical dependencies. This tensor is then normalized via min-max scaling to the [0, 1] interval to enable fair comparison across dimensions.
The normalized MI scores are then structured into a three-dimensional tensor (emotion × frequency band × channel) and visualized topographically using the standard 10–20 system for electrode positioning. A 3×5 grid layout illustrates MI distributions across the Delta (1–4 Hz), Theta (4–8 Hz), Alpha (8–13 Hz), Beta (13–30 Hz), and Gamma (30–50 Hz) bands for each emotional category. Color gradients denote the strength of feature importance, with electrode positions highlighted as black dots. This methodology effectively integrates information-theoretic metrics with neuroimaging visualization, offering intuitive insights into the spatial and spectral dynamics of emotion processing.
TABLE VII: Performance of the SDC-Net model in the ablation study
Method SEED SEED_IV
@@ V-F Topographic Analysis of Important EEG Patterns
V-G Negative Transfer Results
This study evaluates the effectiveness of the proposed SDC-Net framework in alleviating negative transfer in fully unsupervised cross-subject EEG emotion recognition across 45 experimental tasks. Negative transfer is defined as a classification accuracy lower than 33.3% on the SEED dataset or 25% on the SEED_IV dataset. As shown in Table VII, SDC-Net achieved zero instances of negative transfer, demonstrating a clear advantage over conventional transfer learning approaches.
The robustness of SDC-Net in avoiding negative transfer can be attributed to three key components:(1)SS-Mix, which augments intra-subject data while preserving individual-specific characteristics, reducing ambiguity between subject identity and emotion labels;(2) Dynamic Distribution Alignment in RKHS, which jointly aligns marginal and class-conditional distributions using a unified kernel mean embedding framework, while adaptively filtering high-confidence pseudo-labels to ensure alignment quality;(3) DSCL, which imposes structural constraints on pairwise similarities across domains, promoting semantic consistency without relying on temporal alignment. Together, these innovations enable SDC-Net to dynamically adjust to distribution shifts, suppress noisy pseudo-labels, and maintain semantic integrity across subjects, thereby ensuring stable and reliable transfer performance in complex, label-free EEG emotion recognition scenarios.
V-H Why SDC-Net Works? Effectiveness Analysis
@@
Addressing the significant inter-subject variability inherent in EEG data remains a central challenge in the development of robust and generalizable aBCI systems. In this work, we proposed SDC-Net, a novel semantic-dynamic consistency network, and conducted comprehensive evaluations on three publicly available affective EEG datasets: SEED, SEED-IV, and Faced. Experimental results consistently demonstrate that SDC-Net outperforms existing state-of-the-art methods in cross-subject emotion recognition tasks, validating its effectiveness and broad applicability.
A key innovation lies in the SS-Mix module, which generates augmented samples through intra-trial interpolation. Unlike existing GAN-based augmentation methods such as GANSER[19] and SA-cWGAN[20], which risk introducing low-quality or identity-inconsistent samples, SS-Mix explicitly preserves subject identity by avoiding inter-subject mixing. This design enhances intra-subject emotional variability while reducing semantic noise, thus improving model robustness and generalizability. To address the evolving nature of feature distributions, we introduce a dynamic distribution alignment strategy in RKHS, which jointly aligns both marginal and conditional distributions through multi-kernel MMD-based losses. Compared to fixed or marginal alignment approaches—such as BiHDM[34], MS-MDA[7], MS-FRAN[8], DA-CapsNet[35], SDDA[11], PR-PL[13] and DAPLP[10], our method offers finer-grained adaptation to inter-subject and inter-class shifts. Empirical results show that this component alone contributes to a 5.6% accuracy gain and reduces class-level feature overlap, as visualized in the t-SNE plots. Moreover, we propose a novel DSCL mechanism to overcome the limitations of contrastive learning approaches such as CLISA[15] and CL-CS[17], which rely on synchronized stimulus assumptions and suffer from poor generalization under asynchronous conditions. SDC-Net instead infers latent structural similarities within and across domains and selectively enforces consistency constraints without relying on time alignment or pseudo-labels. This enables robust learning of semantic boundaries, even in unlabeled target domains. Ablation experiments confirm that removing this module causes a 3–4% drop in performance, demonstrating its critical role in enhancing discriminability and transferability.
The ablation studies confirm that removing either the structure-aware consistency module or the dynamic reweighting mechanism leads to a 4%-6% performance degradation, highlighting the necessity of each component. Sensitivity analyses (Fig.78) further show that the model maintains high robustness under a wide range of kernel numbers and similarity thresholds, with accuracy fluctuation consistently under 2%. Finally, the robustness of SDC-Net is further validated through visualization and sensitivity analyses. As shown in the t-SNE plots (Fig. 6), emotion categories form more compact clusters with reduced domain discrepancy. Moreover, hyperparameter studies (Fig.7-9) indicate that our method delivers consistently stable performance across a wide range of settings. To further highlight its reliability, we also investigate the risk of negative transfer-a known challenge in domain adaptation where transferred knowledge may impair target performance-and show that SDC-Net is highly resilient to such effects.
In the Future, several promising directions is deserved to research. Firstly, improving pseudo-label quality: One possible enhancement is adjusting the pseudo-label confidence threshold during the later stages of model training to ensure higher label quality. However, setting an excessively high threshold may discard too many uncertain pseudo-labels, leading to an imbalance between categories and low pseudo-label utilization[62, 63]. Therefore, future work should focus on optimizing the balance between the quantity and quality of pseudo-labels, potentially through adaptive thresholds or other robust filtering techniques. Another important avenue is the real-time deployment of the SDC-Net framework within an online aBCI system, facilitating dynamic user interaction. This would allow continuous model updates based on real-time EEG data, facilitating high-precision emotion recognition in dynamic, real-world settings. Testing and integrating our model into such an interactive online aBCI environment would provide a strong foundation for personalized emotion recognition systems that can adapt in real-time.
VI Conclusion
@@
This paper proposes a novel domain adaptation framework to address the challenge of individual variability in aBCIs for cross-subject EEG emotion recognition. The framework integrates three key innovations: SS-Mix data augmentation, dynamic distribution alignment in RKHS, and dual-domain similarity consistency learning Strategy. These components collectively enhance the generalization ability and robustness of emotion recognition models across different subjects.
Experimental results on the SEED, SEED-IV and FACED datasets demonstrate the superiority of the proposed framework in both cross-subject and cross-session scenarios, significantly improving emotion recognition performance. In future work, we plan to explore the integration of online transfer learning techniques to address real-world variability and enable real-time adaptation for practical applications.

### v2

@@ IV-B Confusion matrices
To further analyze classification behavior, we compare the confusion matrices of four representative models: DA-CapsNet, PLMSDANet, LGDAAN-Net, and our proposed SDC-Net. The detailed results are provided in Figure S1 of the Supplementary Information (Section S3.2). The confusion matrices provide insight into how well each model classifies inputs into three categories: Negative, Neutral, and Positive. The diagonal elements represent correct predictions, while off-diagonal elements correspond to misclassifications. Our proposed SDC-Net model outperforms the other models, particularly in classifying the Neutral class with an accuracy of 92.45%, which is the highest among all models.
@@
Additionally, SDC-Net achieves 95.7% accuracy in the Positive class and 87.23% in the Negative class. The confusion between classes is minimal, with only 5.25% of neutral instances being misclassified as negative and 2.53% of positive instances being misclassified as negative. The low misclassification rates in SDC-Net suggest that our model effectively captures subtle differences between sentiment classes, particularly between negative and neutral sentiments, where other models faltered. It achieves the best accuracy in both the Neutral and Positive classes and exhibits significantly lower misclassification rates compared to the other models. This indicates that SDC-Net provides a more nuanced understanding of sentiment, making it particularly effective for tasks that require fine-grained sentiment analysis. In contrast, models such as LGDAAN-Net, while demonstrating strong performance in classifying positive samples, exhibit notable deficiencies in differentiating between negative and neutral sentiments. This limitation significantly undermines their effectiveness in comprehensive sentiment classification tasks.
IV-C Ablation Study
@@
To evaluate the contribution of each module in the proposed SDC-Net, we performed ablation experiments by systematically removing six key components. Table V summarizes the performance under each setting in terms of average classification accuracy. First, domain alignment mechanisms were found to be essential. Removing of the MMD component resulted in a significant drop in performance (86.82%), indicating its crucial role in reducing marginal distribution discrepancy between source and target domains. Similarly, excluding CMMD led to a moderate performance decline (89.91%), showing that conditional alignment further refines domain adaptation, though its impact is secondary to MMD. Second, the effect of DSCL was also prominent. Supervised similarity consistency on source domain contributes to extracting discriminative features from labeled source data, and its removal decreased the accuracy to 88.99%. More critically, excluding self-inferred similarity consistency leanring on target domain reduced performance to 86.99%, emphasizing its importance in modeling intra-class similarity in the unlabeled target domain, thus supporting better generalization. Third, SS-Mix based data augmentation improved generalization by synthesizing diverse EEG trials. Its removal caused a noticeable decrease in Acc (87.99%), suggesting its role in mitigating overfitting and increasing sample diversity. Finally, the pseudo-confidence mechanism, which filters unreliable pseudo-pairs, slightly improved the mean accuracy and significantly reduced variance. Without this mechanism, the model still achieved 89.66%, but with increased performance fluctuation (standard deviation 7.18), indicating its stabilizing effect on target domain predictions.
TABLE V: Performance of the SDC-Net model in the ablation study
Ablation Experiment Strategy Acc (%)
without-$S$-Mix 87.99 $\pm$ 5.77
@@
Figure 3: t-SNE visualization of source and target domain emotion representations for three subjects across training epochs.
@@
Figure 4: Topographic analysis of the mutual information between EEG frequency-band features and model predictions across emotional states.
IV-D Visualization of Domain Alignment via t-SNE
To qualitatively assess the effectiveness of domain alignment, we visualized the feature distributions of source and target domains using t-SNE at different training stages (Fig.3). In these plots, the source domain ($D_{s}$) is represented by circles and the target domain ($D_{t}$) by stars, with emotion categories indicated by color (blue: negative, red: neutral, green: positive). At the early training stage, features from $D_{s}$ and $D_{t}$ are poorly aligned. The target samples, especially for Subject 1 and Subject 2, are scattered and show significant overlap across emotion classes, indicating large domain discrepancy. Positive target samples (green stars) frequently overlap with other categories, suggesting the initial failure in feature alignment. As training progresses, domain alignment improves notably. For Subject 2 and 3, the target domain emotion clusters begin to align with those of the source domain. Notably, negative and positive emotions become more separable, suggesting that MMD and CMMD modules effectively minimize marginal and conditional distribution gaps. By the final training stage, most emotion clusters exhibit clear separation in both domains. Subject 3 shows the most consistent alignment across categories, while Subject 1 retains partial overlap in neutral emotions. This aligns with known challenges in classifying neutral emotional states, which often lie closer to the decision boundaries due to their ambiguous EEG patterns.
IV-E Negative Transfer Results
This study evaluates the effectiveness of the proposed SDC-Net framework in alleviating negative transfer in fully unsupervised cross-subject EEG emotion recognition across 45 experimental tasks. Negative transfer is defined as a classification accuracy lower than 33.3% on the SEED dataset or 25% on the SEED-IV dataset. As shown in TableVI, SDC-Net achieved zero instances of negative transfer, demonstrating a clear advantage over conventional transfer learning approaches.
The robustness of SDC-Net in avoiding negative transfer can be attributed to three key components:(1)SS-Mix, which augments intra-subject data while preserving individual-specific characteristics, reducing ambiguity between subject identity and emotion labels;(2) Dynamic Distribution Alignment in RKHS, which jointly aligns marginal and class-conditional distributions using a unified kernel mean embedding framework, while adaptively filtering high-confidence pseudo-labels to ensure alignment quality;(3) DSCL, which imposes structural constraints on pairwise similarities across domains, promoting semantic consistency without relying on temporal alignment. To sum up, these innovations enable SDC-Net to dynamically adjust to distribution shifts, suppress noisy pseudo-labels, and maintain semantic integrity across subjects, thereby ensuring stable and reliable transfer performance in complex, label-free EEG emotion recognition scenarios.
TABLE VI: Number of Subjects Exhibiting Negative Transfer on SEED and SEED-IV Datasets
Method SEED SEED-IV
@@ IV-E Negative Transfer Results
IV-F Topographic Analysis of Important EEG Patterns
This study utilizes an EEG topographic mapping approach grounded in Mutual Information (MI) to evaluate the contributions of different brain regions and frequency bands to emotion classification. The preprocessed EEG feature matrix $\mathrm{\mathbf{X}}\in\mathbb{R}^{N\times 5\times 2}$, incorporating signals from 62 electrode channels across five frequency bands, is paired with the classification probability matrix $\mathrm{\mathbf{Y}}\in\mathbb{R}^{N\times 3}$, corresponding to three emotional states: negative, neutral, and positive. By computing the MI between each EEG feature and the predicted probabilities of the emotion categories, an initial MI $MI\in R^{3\times 5\times 2}$ tensor is derived, capturing the nonlinear statistical dependencies. This tensor is then normalized via min-max scaling to the [0, 1] interval to enable fair comparison across dimensions. The normalized MI scores are then structured into a three-dimensional tensor (emotion × frequency band × channel) and visualized topographically using the standard 10–20 system for electrode positioning. A 3×5 grid layout illustrates MI distributions across the Delta (1–4 Hz), Theta (4–8 Hz), Alpha (8–13 Hz), Beta (13–30 Hz), and Gamma (30–50 Hz) bands for each emotional category. Color gradients denote the strength of feature importance, with electrode positions highlighted as black dots. This methodology effectively integrates information-theoretic metrics with neuroimaging visualization, offering intuitive insights into the spatial and spectral dynamics of emotion processing.
IV-G Discussion
@@
Addressing significant inter-subject variability in EEG data remains a central challenge in developing robust and generalizable aBCI systems. Beyond accurate classification, the ultimate goal of aBCIs is to achieve stable, interpretable, and scalable emotion measurement across users and scenarios. To this end, we propose SDC-Net, designed to improve not only recognition performance but also the consistency and robustness of emotion measurement. We conducted comprehensive evaluations on three public EEG emotion datasets-SEED, SEED-IV, and FACED-demonstrating that SDC-Net significantly outperforms state-of-the-art methods in cross-subject emotion recognition tasks, validating its effectiveness and broad applicability in affective computing and intelligent instrumentation.
One of the key innovations in SDC-Net is the SS-Mix module, which generates augmented samples through intra-trial interpolation. Compared with GAN-based data augmentation methods (e.g., GANSER[21] and SA-cWGAN[22]), which may introduce low-quality or identity-inconsistent samples, SS-Mix enhances the modeling of intra-subject emotional variability while significantly reducing semantic noise. From a measurement perspective, this approach can be regarded as a strategy for increasing reliable sample density without introducing cross-subject artifacts, which is especially important in single-trial settings or when data is limited. To address the limitations of prior methods that rely on static distribution alignment and fixed pseudo-labels-such as BiHDM[29], MS-MDA[9], MS-FRAN[10], DA-CapsNet[30], SDDA[13], PR-PL[15], and DAPLP[12]-we propose DDA strategy in RKHS, which provides finer adaptation to inter-subject and inter-class distributional shifts. This strategy also incorporates a confidence-based pseudo-label filtering mechanism that dynamically selects target samples with well-defined semantic structures, enabling progressive alignment from easy to hard examples. Functionally, this design can be viewed as a robust calibration mechanism when labels are noisy or incomplete, significantly improving training stability and generalization. The proposed DSCL module further addresses the limitations of existing contrastive learning methods (e.g., CLISA[17], CL-CS[19]) that rely on temporal synchronization and are difficult to generalize to cross-scenario measurement. DSCL infers latent structural similarity between domains and imposes corresponding consistency constraints, enabling semantic boundary learning without time alignment or pseudo-label supervision. Experimental results suggest DSCL functions as latent structural regularization, guiding the model toward learning compact and well-separated topological structures.
Ablation studies indicate that removing either the structural consistency module or the dynamic reweighting mechanism results in a 4%–6% drop in performance, highlighting their critical roles. Sensitivity analysis further shows that SDC-Net maintains high robustness under different kernel numbers and similarity thresholds, with performance fluctuation consistently below 2% (detailed results are provided in Supplementary information (Section S4)). The robustness of SDC-Net is also validated through visualization and negative transfer testing. As shown in Fig. 3, emotional categories form well-defined and separable clusters across source and target domains. Additionally, SDC-Net demonstrates strong resistance to negative transfer, ensuring reliable emotion measurement in cross-subject settings.
V Conclusion
@@
This paper proposes SDC-Net, a novel domain adaptation framework designed to address the challenge of individual variability in aBCIs for cross-subject EEG emotion recognition. The framework integrates three key innovations: (1) the Same-Subject Same-Trial Mixup data augmentation strategy; (2) dynamic distribution alignment in the RKHS; and (3) dual-domain similarity consistency learning strategy. Collectively, these components significantly enhance the capability of generalization and robustness for emotion recognition models across different subjects. Extensive experimental results on the SEED, SEED-IV, and FACED datasets demonstrate the superiority of SDC-Net over existing methods in both cross-subject and cross-session scenarios, achieving significant improvements in emotion recognition performance.
Future work may focus on two promising directions. First, enhancing pseudo-label quality by balancing label confidence and quantity remains crucial. Adaptive confidence thresholds or robust filtering strategies could help mitigate the trade-off between discarding noisy labels and preserving class diversity [51, 52]. Second, deploying the SDC-Net framework in real-time aBCI systems represents a crucial step toward dynamic, user-adaptive emotion measurement. Integration with online EEG acquisition platforms could enable continuous learning from streaming signals, facilitating personalized and context-aware affective computing in real-world scenarios, such as clinical monitoring or closed-loop neurofeedback instrumentation.